OpenStudy (anonymous):

What is the easiest way to solve this: 5x^2-180=0?

pooja195 (pooja195):

Hi welcome to openstudy :) You can add 180 to both sides then use the differences of 2 squares method

pooja195 (pooja195):

\[\huge~\rm~5x^2−180+180=0+180\]

pooja195 (pooja195):

Can you tell me what equation we would have once we add 180 to both sides?

OpenStudy (anonymous):

5x^2=180

pooja195 (pooja195):

Good now let's try isolating x 5/5=x^2 180/5= ?

OpenStudy (anonymous):

180/5=36

pooja195 (pooja195):

Good Now we have \[\huge~\rm~x^2=36\] Let's use the difference of 2 squares method basically all we do is square root both numbers :) \[\huge~\rm~\sqrt{x^2}=\sqrt{36}\] Can you tell me what the answer is? :)

OpenStudy (anonymous):

x=6

pooja195 (pooja195):

Correct but we would have \[\huge~\rm~x=6 ~AND~ x=-6\]because we square rooted the 36 there are 2 possible answers :)

OpenStudy (anonymous):

thank you:)

pooja195 (pooja195):

yw :D ^_^ \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)